OpenStudy (anonymous):

Indicate in standard form the equation of the line passing through the given points. G(4, 6), H(1, 5)

OpenStudy (jdoe0001):

\(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) G&({\color{red}{ 4}}\quad ,&{\color{blue}{ 6}})\quad % (c,d) H&({\color{red}{ 1}}\quad ,&{\color{blue}{ 5}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{rise}{run} \implies \cfrac{{\color{blue}{ y_2}}-{\color{blue}{ y_1}}}{{\color{red}{ x_2}}-{\color{red}{ x_1}}} \\ \quad \\ % point-slope intercept y-{\color{blue}{ y_1}}={\color{green}{ m}}(x-{\color{red}{ x_1}})\qquad \textit{plug in the values and move everything to the left}\\ \qquad \uparrow\\ \textit{point-slope form}\)

OpenStudy (texaschic101):

first we find the slope by using the slope formula slope(m) = (y2 - y1) / (x2 - x1) (4,6)...x1 = 4 and y1 = 6 (1,5)...x2 = 1 and y2 = 5 now we sub slope(m) = (5 - 6) / (1 - 4) can you finish and find the slope ?

OpenStudy (anonymous):

-1/-3?

OpenStudy (jdoe0001):

yeap, the slope is\(\bf \cfrac{-1}{-3}\implies \cfrac{1}{3}\) thus now plug in the \(x_1 \ and \ y_1\) values in the point-slope form \(\bf \begin{array}{lllll} &x_1&y_1&x_2&y_2\\ % (a,b) G&({\color{red}{ 4}}\quad ,&{\color{blue}{ 6}})\quad % (c,d) H&({\color{red}{ 1}}\quad ,&{\color{blue}{ 5}}) \end{array} \\\quad \\ % slope = m slope = {\color{green}{ m}}= \cfrac{1}{3} \\ \quad \\ % point-slope intercept y-{\color{blue}{ 6}}={\color{green}{ \cfrac{1}{3}}}(x-{\color{red}{ 4}})\qquad \textit{plug in the values and move everything to the left}\\ \qquad \uparrow\\ \textit{point-slope form}\) so, when you move everything to the left-side, the only thing left on the right-side will be a zero =0